Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2019-01-29

Observers: Silvia Piranomonte, Roberta Carini
Support Team: S. Benetti, L. Tomasella, A. Fiore
TAT: Janet Chen, Matt Nicholl, Steve Williams


Calibrations: 
EFOSC2:
biases
flat, arcs gr11, gr13, gr16 (evening )

Weather: 
Start: Humidity 79%, 
Middle: (5:40 UT ) Humidity 58%, we start to observe, the seeing is very bad, there are some technical problems


Comments: 
00:30 ut: the humidity is too high, we are closed
05:40 ut: the TO opens the dome, we can observe, but there are some technical problems.the seeing is about 2.2"

Instrument - EFOCS

UT            Target               Type                Setup                  airmass      Seeing       texp         notes
06:05        LTT3864           STD    gr11,gr16,gr13,slit1,1.5        1.01             1.7"                 
06:27       SN2019jf           follow-up          gr13,sl 1.5               1.20             1.6"      1x1800s
07:10       GRB190129B                             r,R                          1.75             1.8"      2x120s        grb is not visible
07:21       GRB190129B                             r                             1.85             1.8"       2x240s       
07:45       SN2019so         follow-up           g11 g16, sl1.5         1.04            1.8"       2x1800s      at the end seeing 1.08"
09:11        LTT3864           STD                gr11,gr16,gr13,1.5    1.2               1.08     

Comments